# Sliding Puzzles Gym: A Scalable Benchmark for State Representation in Visual Reinforcement Learning

### v4

@@ 1 Introduction
While traditional RL benchmarks such as Atari (Bellemare et al., 2013) and DeepMind Control Suite (Tassa et al., 2018) provide valuable testing grounds for overall agent performance, they inherently intertwine representation learning with policy optimization and environment dynamics. More recent specialized benchmarks for visual learning, though promising, fall short in their ability to precisely modulate visual complexity. ProcGen (Cobbe et al., 2020) concurrently alters both visual and task difficulty, obscuring the specific impact of representation learning, while the Distracting Control Suite (Stone et al., 2021) incorporates visual variations that are orthogonal to the core task objectives. Consequently, existing benchmarks lack the necessary precision to systematically assess an agent’s capacity to acquire task-relevant visual representations in isolation.
@@
To address this critical gap in visual RL evaluation, we introduce the Sliding Puzzles Gym (SPGym) [^1] Available at https://github.com/bryanoliveira/sliding-puzzles-gym., an open-source benchmark specifically designed to assess how agents handle increasingly diverse visual observations. As depicted in Figure 1, SPGym transforms the classic 8-tile puzzle into a visual RL task through three innovative design principles: (1) maintaining consistent environment dynamics regardless of difficulty level, ensuring that the underlying task remains fixed; (2) enabling precise control over visual complexity through adjustable grid dimensions and image pool sizes; and (3) establishing an unambiguous success metric based on puzzle completion. This carefully engineered framework provides researchers with a controlled environment to systematically evaluate how agents’ performance degrades with increasing visual diversity, offering insights into the fundamental limitations of their representation learning capabilities while keeping the core task unchanged.
Our experiments with SPGym reveal that it effectively distinguishes agents by their representation learning capabilities, showing a strong correlation with task performance. While pretrained encoders and data augmentation provide clear benefits, many advanced methods surprisingly underperform their baselines, suggesting their assumptions don’t transfer to SPGym’s unique blend of visual complexity and structured dynamics. More concerningly, our tiered generalization analysis exposes fundamental limitations: agents achieve perfect performance on training images but fail to transfer to unseen ones, even with larger training pools. Performance often degrades as pool size increases, highlighting a key trade-off: greater visual diversity increases learning difficulty without improving generalization, suggesting current methods struggle to learn truly generalizable representations. We further demonstrate SPGym’s extensibility through experiments with procedurally generated images and larger puzzles, maintaining fixed observation/action spaces while expanding state space and visual diversity. These findings expose critical gaps in current visual RL methods and establish SPGym as a valuable tool for advancing robust, generalizable decision-making systems.
@@
We make three key contributions to visual RL research: (1) SPGym, a novel benchmark that systematically evaluates representation learning by scaling visual complexity while keeping environment dynamics constant; (2) an extensive empirical analysis of state-of-the-art methods, uncovering critical limitations in their ability to process diverse visual inputs; and (3) fundamental insights into the challenges of scaling visual RL, offering directions for advancing representation learning in decision-making systems.
@@ 2 Related Work
Foundational visual RL benchmarks like the Atari Learning Environment (Bellemare et al., 2013), DeepMind Control Suite (Tassa et al., 2018), DeepMind Lab (Beattie et al., 2016), and CARLA (Dosovitskiy et al., 2017) pioneered pixel-based agent evaluation. These environments drove advances in representation learning techniques including pretraining (Higgins et al., 2017; Stooke et al., 2021; Schwarzer et al., 2021), contrastive learning (Laskin et al., 2020b), self-supervised prediction (Schwarzer et al., 2020), and world models (Hafner et al., 2025). However, by assessing overall agent performance across representation learning, policy optimization, and dynamics modeling, these benchmarks obscure the specific impact of representation learning improvements.
SPGym addresses this limitation by precisely controlling visual complexity while keeping environment dynamics fixed, enabling focused evaluation of representation learning.
@@
Recent benchmarks have sought to better assess visual generalization. ProcGen (Cobbe et al., 2020) employs procedurally generated levels, but its difficulty scaling simultaneously affects both visual and task complexity. The Distracting Control Suite (Stone et al., 2021) evaluates robustness to visual variations, but these serve as task-irrelevant distractions that agents can safely ignore while solving core tasks. COOM (Tomilin et al., 2023) provides a continual learning benchmark for embodied pixel-based RL in 3D environments, focusing on catastrophic forgetting and knowledge transfer across tasks rather than isolating representation learning in a single task. These approaches fail to create representation learning challenges essential for task success.
SPGym overcomes these limitations by making visual understanding fundamental to puzzle solving while maintaining fixed task complexity. Through controlled scaling of visual diversity in puzzle tiles, SPGym creates representation learning challenges tied to task completion, enabling precise evaluation of agents’ visual learning capabilities.
@@
Estermann et al. (2024) demonstrated puzzle-based environments’ value for evaluating neural algorithmic reasoning, focusing on discrete states. Though they tested pixel observations with sparse rewards, agents struggled with basic visual inputs. This work revealed puzzles’ potential and challenges for visual learning, showing the need for systematic representation learning evaluation in these controlled settings.
SPGym advances this research direction by incorporating rich visual observations while maintaining the controlled nature of puzzle environments, facilitating systematic assessment of representation learning in the visual domain.
@@
Classical approaches to sliding puzzles employ heuristic search algorithms like A* and IDA* with domain-specific heuristics such as Manhattan distance (Korf, 1985; Burns et al., 2012; Lee & See, 2022). While these methods guarantee optimal solutions, they require direct access to internal puzzle states and are computationally demanding. Deep RL provides a scalable alternative that learns effective strategies without hand-crafted heuristics (Agostinelli et al., 2019; Moon & Cho, 2024; Estermann et al., 2024). However, prior work has predominantly focused on discrete state representations rather than learning from visual observations (Agostinelli et al., 2019; Moon & Cho, 2024).
Building on this foundation, our work evaluates standard RL algorithms on SPGym, where agents must learn exclusively from pixel observations without access to internal states, offering insights into their out-of-the-box performance in this challenging visual learning setting.
@@ 3 The Sliding Puzzles Gym
SPGym extends the classic sliding puzzle, a game where players rearrange shuffled tiles on a grid by sliding a tile into an adjacent empty space, with the goal of restoring an ordered configuration (Figure 1). Our framework generalizes this by supporting configurable $H\times W$ grid dimensions (from $2\times 2$ upwards) and diverse observation modalities. In our experiments, we primarily use a $3\times 3$ grid where tiles are image patches, and the agent receives the composite image of the grid as input. SPGym adheres to the Gym (Brockman, 2016) interface, promoting modularity between environment and agent.
@@
We formulate SPGym as a partially observable Markov decision process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{X},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{S}_{0})$, where $\mathcal{S}$ is the finite state space representing the internal configuration of puzzle tiles, $\mathcal{X}:$ is the observation space of image-based tiles, $\mathcal{A}$ is the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ defines the deterministic transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, and $\mathcal{S}_{0}$ is the initial state distribution.
@@
Crucially, agents do not have direct access to states $s\in\mathcal{S}$ and must learn policies $\pi:\mathcal{X}\rightarrow\mathcal{A}$ based solely on observations $x\in\mathcal{X}$. For visual observations, each training run begins by sampling $p$ images from a predefined dataset to form an image pool $\mathcal{I}$. At the start of each episode, we select a random image $i\in\mathcal{I}$ and split it into $H\times W$ distinct, spatially indexed patches, and overlay it onto the puzzle. The agent’s task is therefore to reconstruct the shuffled image, testing its ability to form compositional representations from pixels. This formulation provides two independent mechanisms for controlling complexity: (1) varying $p$ adjusts the diversity of the observation space $\mathcal{X}$ by changing the pool of available images, and (2) modifying $H$ and $W$ alters the state space complexity by changing the grid dimensions and number of puzzle pieces. Both mechanisms operate while keeping the underlying dynamics $\mathcal{P}$, the action space $\mathcal{A}$, and the reward function $\mathcal{R}$ fixed. SPGym is dataset-agnostic, supporting any image dataset, including procedurally generated ones.
@@
The action space $\mathcal{A}$ contains four possible actions: UP, DOWN, LEFT, or RIGHT, which move the corresponding tile to the empty space in a discrete manner. Once the agent selects a tile to move, the dynamics $\mathcal{P}$ define the next puzzle state in a predictable and deterministic way, as illustrated in Figure 1. The objective is to rearrange shuffled image tiles into their correct positions, forming a complete, intelligible image.
@@
The reward function $\mathcal{R}$ provides dense feedback $r$ to guide agent learning while accommodating the non-monotonic nature of optimal puzzle-solving paths. Following established approaches in the sliding puzzle literature (Korf, 1985; Burns et al., 2012; Lee & See, 2022; Moon & Cho, 2024), we base our reward on the Manhattan distance between each tile’s current position and its target position, normalized across all tiles. This distance metric is widely adopted in computational solutions to sliding puzzles due to its efficiency and effectiveness in capturing progress toward the goal state. Specifically, at each time step, the reward is computed as follows:
$r_{t}=\begin{cases}-D,&\text{if action is valid}\\ {\color[rgb]{0.55,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.55,0.0,0.0}-1,}&\text{if action is invalid}\\ {\color[rgb]{0.0,0.39,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.39,0.0}+1,}&\text{if puzzle is solved}\end{cases}\text{\hskip 10.00002pt, with}$ (1)
@@
Here, $(u_{i,j},v_{i,j})$ represents the current position of the tile at index $(i,j)$, $(u^{*}_{i,j},v^{*}_{i,j})$ is its target position. $D$ is the normalized Manhattan distance between current and target positions. Invalid actions, such as attempting to move a tile outside the grid boundaries, result in a penalty of ${\color[rgb]{0.55,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.55,0.0,0.0}-1}$ and don’t alter the puzzle state. For example, in Figure 1, the DOWN and RIGHT actions would be invalid for the first state. Successfully solving the puzzle rewards the agent with ${\color[rgb]{0.0,0.39,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.39,0.0}+1}$ and terminates the episode. This formulation provides a well-behaved learning signal between $[-1,+1]$ and encourages solving the puzzle in minimal steps.
@@
While our main experiments focus on $3\times 3$ grids with image observations, SPGym supports variable grid sizes from $2\times 2$ to larger $H\times W$ configurations. We demonstrate this with experiments on $4\times 4$ grids (Table 3), which significantly increase sample complexity compared to $3\times 3$ grids. We also use one-hot observation spaces, illustrated in Figure 2, as a proxy for ground-truth puzzle states to benchmark the visual representation challenge (results in Appendix C.1).
@@
The primary scaling mechanism in SPGym involves increasing visual diversity by expanding the image pool while keeping the grid size fixed (Figure 3). As shown in Table 1, larger image pools consistently increase sample complexity. Since core task elements like state and action spaces, transition dynamics, and reward function remain unchanged, this increased difficulty primarily reflects the challenge of learning robust visual representations. The quality of learned representations is directly linked to task performance, as demonstrated by linear probe analyses (Appendix C.1.2). Probe accuracy decreases with larger pool sizes (Table 10), exhibiting strong correlation with sample efficiency. Comparisons with one-hot encoded baselines (Appendix C.1.1), where the representation learning burden is minimal, further highlight this focus. While the one-hot setting offers fixed difficulty, the systematic increase in sample complexity for image-based agents as visual diversity scales provides a stress test not possible in the one-hot setting.
@@
The secondary scaling mechanism adjusts grid dimensions. Increasing from $3\times 3$ to $4\times 4$ dramatically increases the number of steps to solve the task due to expanded state space complexity. While $3\times 3$ puzzles have approximately $1.81\times 10^{5}$ possible states, $4\times 4$ puzzles present a dramatically larger challenge with approximately $10^{13}$ possible states. This increase in puzzle size not only expands the state space exponentially but also makes the representation learning problem harder, as the same image must now be divided into more patches that need to be correctly positioned. Importantly, larger grids don’t affect the observation space $\mathcal{X}$ (image sizes remain fixed) or action space $\mathcal{A}$ (still four directional moves), and both puzzle sizes maintain identical observation and action spaces, isolating the impact of state space complexity. While this mechanism evaluates exploration and policy-learning capabilities, our experiments demonstrate that $3\times 3$ grids provide sufficient discriminative power for visual representation learning assessment.

### v5

@@ 1 Introduction
While traditional RL benchmarks such as Atari (Bellemare et al., 2013) and DeepMind Control Suite (Tassa et al., 2018) offer valuable evaluation platforms for overall agent effectiveness, their performance metrics inherently intertwine representation learning with policy optimization and environment dynamics. More recent dedicated benchmarks for visual learning, though promising, fall short in their capacity to precisely modulate visual complexity. For instance, ProcGen (Cobbe et al., 2020) simultaneously modifies both visual and task difficulty, obscuring the specific impact of representation learning, while the Distracting Control Suite (Stone et al., 2021) introduces visual distractors that are unrelated to the main task and can be safely disregarded by the model. Consequently, current benchmarks lack the necessary precision to systematically assess an agent’s ability to acquire task-relevant visual representations.
@@
To address this gap, we introduce the Sliding Puzzles Gym (SPGym) [^1] Available at https://github.com/bryanoliveira/sliding-puzzles-gym., an open-source benchmark created to evaluate how agents handle increasingly diverse visual observations across training runs. As depicted in Figure 1, SPGym transforms the classic 8-tile puzzle into a visual RL challenge through three central design principles: (1) preserving consistent environment dynamics regardless of difficulty, ensuring that the underlying task remains unchanged; (2) enabling precise control over visual complexity through adjustable grid dimensions and image pool sizes; and (3) establishing a clear success metric based on puzzle completion. This carefully engineered framework provides researchers with a controlled environment to systematically examine how agent performance degrades as visual diversity increases, offering insights into the fundamental limitations of their representational abilities while learning from experience.
Our experiments show that SPGym effectively distinguishes agents by their out-of-the-box representation learning abilities, with strong correlations between linear probing accuracy of learned encoders and task performance. While pretraining and data augmentation are beneficial, many advanced methods with standard configurations for discrete visual RL underperform, suggesting either a need for task-specific tuning or fundamental mismatches with SPGym’s visual-structural dynamics. More concerningly, our tiered generalization analysis exposes fundamental limitations: on SPGym, agents that master the training images fail to transfer to unseen ones, even when trained on larger and more diverse image pools. In fact, performance often degrades as visual diversity increases, indicating that current methods struggle to learn truly generalizable representations. We further demonstrate SPGym’s extensibility through experiments with procedurally generated images and larger puzzles, maintaining fixed observation/action spaces while expanding state space and visual diversity. These findings expose critical gaps in current visual RL methods and establish SPGym as a valuable tool for advancing robust, generalizable artificial agents.
@@
We make three key contributions to visual RL research: (1) we introduce SPGym, a novel benchmark that systematically evaluates representation learning by scaling visual complexity while keeping environment dynamics constant; (2) we conduct an extensive empirical analysis of state-of-the-art methods, uncovering critical limitations in their ability to process diverse visual inputs; and (3) we provide fundamental insights into the challenges of scaling visual RL, offering directions for advancing representation learning in decision-making systems.
@@ 2 Related Work
Visual RL environments such as the Arcade Learning Environment (Bellemare et al., 2013), DeepMind Control Suite (Tassa et al., 2018), DeepMind Lab (Beattie et al., 2016), and CARLA (Dosovitskiy et al., 2017) are widely used for pixel-based agent evaluation. These platforms have driven progress in visual representation learning for decision-making, including pretraining (Higgins et al., 2017; Stooke et al., 2021; Schwarzer et al., 2021), contrastive methods (Laskin et al., 2020b), self-supervised prediction (Schwarzer et al., 2020), and world models (Hafner et al., 2025). However, by measuring overall agent effectiveness, which relies on learning representations, policies, and dynamics, the specific impact of advances in representation learning for RL may be obscured. SPGym addresses this drawback by regulating visual complexity while maintaining constant environment dynamics, allowing for targeted assessment of representation learning approaches for decision-making.
@@
Recent benchmarks have sought to better assess visual generalization and robustness of RL agents. ProcGen (Cobbe et al., 2020) employs procedurally generated levels, but its difficulty scaling simultaneously affects both visual and task complexity. The Distracting Control Suite (Stone et al., 2021) evaluates robustness to visual variations, but these serve as task-irrelevant distractions that agents can safely learn to ignore while solving core tasks. COOM (Tomilin et al., 2023) provides a continual learning benchmark for embodied pixel-based RL in 3D environments, focusing on catastrophic forgetting and knowledge transfer across tasks rather than isolating representation learning in a single task. However, these approaches do not make visual understanding essential for task success. SPGym overcomes this limitation by making visual coherency fundamental to puzzle solving while maintaining fixed task complexity. Through controlled scaling of visual diversity in puzzle tiles, SPGym ensures that the representation learning challenges are tied to task completion, enabling a more precise evaluation of RL agents’ visual learning capabilities.
@@
Estermann et al. (2024) demonstrated puzzle-based environments’ value for evaluating neural algorithmic reasoning, focusing on discrete state spaces. Though they tested pixel observations with sparse rewards, agents struggled with basic visual inputs. This work revealed puzzles’ potential and challenges for visual learning, showing the need for systematic representation learning evaluation in these controlled settings. SPGym advances this research direction by incorporating rich visual observations while maintaining the controlled nature of puzzle environments, facilitating systematic assessment of representation learning in the visual domain.
@@
Classical approaches to sliding puzzles employ search algorithms such as A* and IDA*, leveraging domain-specific heuristics like the Manhattan distance (Korf, 1985; Burns et al., 2012; Lee & See, 2022). These methods guarantee optimal solutions but require direct access to the internal puzzle state and are often computationally intensive. Deep RL offers a scalable alternative by learning effective strategies without hand-crafted heuristics (Agostinelli et al., 2019; Moon & Cho, 2024; Estermann et al., 2024), though prior work has primarily focused on discrete state representations rather than learning directly from visual observations (Agostinelli et al., 2019; Moon & Cho, 2024). We build on these methods by using the Manhattan distance as the basis for our reward function, providing a well-shaped learning signal for RL agents. In SPGym, we evaluate standard RL algorithms where agents must learn solely from pixel observations, without access to internal states, to assess their out-of-the-box performance in this challenging setting.
@@ 3 The Sliding Puzzles Gym
SPGym extends the classic sliding tile puzzle, a game where players rearrange shuffled tiles on a grid by sliding a tile into an adjacent empty space, with the goal of restoring an ordered configuration (Figure 1). Our framework generalizes this setup by supporting configurable $H\times W$ grid dimensions (from $2\times 2$ upwards) and diverse observation modalities. In our experiments, we primarily use a $3\times 3$ grid where tiles are image patches, and the agent receives the composite image of the grid as input. Our implementation builds on the Gym (Brockman, 2016) interface for modularity between environment and agent.
@@
We formulate SPGym as a partially observable Markov decision process (POMDP) defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{S}_{0},\Omega,\mathcal{O})$, where $\mathcal{S}$ is the finite state space representing all solvable puzzle configurations, $\mathcal{A}$ is the discrete action space (tile movements), $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ defines the deterministic transition dynamics, $\mathcal{R}:\mathcal{S}\rightarrow\mathbb{R}$ is the reward function, $\mathcal{S}_{0}$ is the initial state distribution (e.g., uniform over $\mathcal{S}$), $\Omega$ is the observation space, and $\mathcal{O}:\mathcal{S}\times\mathcal{I}\rightarrow\Omega$ is the observation function parameterized by a data pool $\mathcal{I}$, which in this work consists of images.
@@
Crucially, agents do not have direct access to states $s\in\mathcal{S}$ and must learn policies based solely on observations in $\Omega$. For visual observations, each training run begins by sampling $p$ images from a predefined dataset to form an image pool $\mathcal{I}$. At the start of each episode, we select a random image $i\in\mathcal{I}$, partition it into $H\times W$ distinct patches corresponding to the puzzle tiles, and assign each patch to a tile position according to the current state $s$. The observation function $\mathcal{O}(s,i)$ then renders the composite image by arranging these patches according to the puzzle configuration. The agent’s task is to reconstruct the original image, testing its ability to form compositional representations from visual input.
This formulation provides two independent mechanisms for controlling complexity: (1) varying the pool size $p$ adjusts the diversity of observations by changing the number of available images in $\mathcal{I}$, directly affecting $|\Omega|$, and (2) modifying grid dimensions $H\times W$ alters both the state space size $|\mathcal{S}|$ and the visual complexity of individual observations. Both mechanisms operate while keeping the underlying transition dynamics $\mathcal{P}$, action space $\mathcal{A}$, and reward function $\mathcal{R}$ invariant. SPGym is dataset-agnostic, supporting any image dataset, including procedurally generated ones.
@@
The action space $\mathcal{A}$ contains four possible actions: UP, DOWN, LEFT, or RIGHT, which move the corresponding tile to the empty space in a discrete manner. Once the agent selects a tile to move, the dynamics $\mathcal{P}$ define the next puzzle state in a predictable and deterministic way, as illustrated in Figure 1.
@@
The reward function $\mathcal{R}$ provides dense feedback to guide agent learning while accommodating the non-monotonic nature of optimal puzzle-solving paths. Following established approaches in the sliding puzzle literature (Korf, 1985; Burns et al., 2012; Lee & See, 2022; Moon & Cho, 2024), we base our reward on the Manhattan distance between each tile’s current position and its target position, normalized across all tiles. This distance metric is widely adopted in computational solutions to sliding puzzles due to its efficiency and effectiveness in capturing progress toward the goal state. Specifically, at each time step, the reward is computed as follows:
[…]
@@
Here, $(u_{i,j},v_{i,j})$ represents the current position of the tile at index $(i,j)$, $(u^{*}_{i,j},v^{*}_{i,j})$ is its target position. $D$ is the normalized Manhattan distance between current and target positions. Invalid actions, such as attempting to move a tile outside the grid boundaries, result in a penalty of $\color[rgb]{0.55,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.55,0.0,0.0}-1$} and don’t alter the puzzle state. For example, in Figure 1, the DOWN and RIGHT actions would be invalid for the first state. Successfully solving the puzzle rewards the agent with $\color[rgb]{0.0,0.39,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.39,0.0}+1$} and terminates the episode. This formulation provides a well-behaved learning signal between [1,+1][-1,+1] and encourages solving the puzzle in minimal steps.
@@
As previously mentioned, SPGym provides two orthogonal mechanisms for scaling task difficulty: visual diversity and grid size. While our primary experiments use $3\times 3$ image-based puzzles, SPGym supports larger grids and other observation modalities, such as one-hot encodings (Figure 2), to isolate specific research questions.
@@
The primary scaling mechanism in SPGym isolates the challenge of representation learning by increasing visual diversity, illustrated in Figure 3, while keeping all other task components fixed across training runs. As the image pool size grows, agents require more samples to solve the task (Table 1), even though the state space, action space, and transition dynamics remain unchanged. By design, this increased difficulty stems directly from the representation learning challenge. We further confirm this link through linear probe analysis (Appendix C.1.2), which shows a strong correlation between representation quality and task performance. This controlled scaling provides a stress test for visual learning that is not possible in settings with fixed visual inputs or state-based observations (see Appendix C.1.1).
@@
The secondary scaling mechanism modulates the search challenge by adjusting grid dimensions. Increasing the grid size from $3\times 3$ to $4\times 4$ dramatically expands the state space complexity and, as a result, the number of steps required to solve the task (see Table 3). For example, $3\times 3$ puzzles have approximately $\frac{3^{2}!}{2}=1.81\times 10^{5}$ possible states, whereas $4\times 4$ puzzles pose a much greater challenge with roughly $\frac{4^{2}!}{2}=10^{13}$ possible states. This increase in grid size also makes representation learning more difficult, since the agent must correctly map and position a greater number of image patches, modeling a larger observation space. However, the size and shape of each rendered observation (i.e., the observation dimensions) remain unchanged, as does the action space (still four directional moves), and the task dynamics continue to be deterministic and fully predictable. While this scaling mechanism enables evaluation of exploration and policy-learning capabilities, our experiments show that $3\times 3$ grids already provide sufficient discriminative power for assessing visual representation learning.